OpenStudy (anonymous):

Please help! 1 f(x)= --- +1 Find x if f(x)=-1. x-2

OpenStudy (anonymous):

Plug in -1 for f(x) and find what x is.

OpenStudy (anonymous):

\[-1 = \frac{1}{x+2} +1\]

OpenStudy (anonymous):

f(-1)= i/(-1-2)+ 1 = -1/3 + 3 / 3 = -2/3.

OpenStudy (anonymous):

That's not right..

OpenStudy (anonymous):

sux, can you solve what I wrote from there?

OpenStudy (anonymous):

I read x - 2 in the denomator, not x + 2.

OpenStudy (anonymous):

you are right, but even so you solved for f(-1) not f(x) = -1.

OpenStudy (anonymous):

\[-1 = \frac{1}{x-2} + 1\] Is what needs solving for x.

OpenStudy (anonymous):

That's so right. My misread :^)

OpenStudy (anonymous):

Np. Now we just need mathsux to finish solving it.

OpenStudy (anonymous):

Easily done.

OpenStudy (anonymous):

For you perhaps, but it's not your problem ;p

OpenStudy (anonymous):

Answer is -2 right?